OpenStudy (anonymous):

A bag has 5 red marbles, 6 blue marbles and 4 black marbles. What is the probability of picking a blue marble, replacing it, and then picking a black marble?

OpenStudy (ganpat):

total marbles in bag = 15.. Number of a blue marble = 6 So probability = 6/15.. Now keeping it in bag and picking black marble = 4/15.. So probability = (6/15) * (4/15)...

OpenStudy (anonymous):

Thank you! that was very helpful!

OpenStudy (ganpat):

great.. so u got the answer.. i assume..